OpenStudy (anonymous):

solve for x 49^(3x+1)=7^((x^2)+16x+18)

OpenStudy (anonymous):

(3x+1)ln(49)=(x^2+16x+18)ln(7) since ln(7)^2=ln(49) using this 2ln(7)=ln(49) 2(3x+1)ln(7)=(x^2+16x+18)ln(7) the ln(7) cancel leaving 6x+2=x^2+16x+18 0=x^2+10x+16 (x+8)(x+2)=0 x=-8 x=-2

OpenStudy (anonymous):

using the rule of ln is this ln(x)^n=nln(x) thats why i brought down everything to the front of the (ln) in case you were curious